# PathoDuet: Foundation Models for Pathological Slide Analysis of H&E and IHC Stains

### v1

@@ 3.2 Cross-scale positioning
When observing how pathologists view slides, the most frequent action is likely to be zooming in and out. This phenomenon originates from the diagnostic process wherein pathologists first detect suspicious regions from a global view, then zoom in and define the region with its local surroundings, and subsequently zoom out to inspect the next region of interest. The underlying concept is that they obtain a global and coarse understanding of a WSI under low magnification, while a local but fine understanding of critical regions under high magnification. Inspired by this, we abstract the concept as a larger region containing a smaller patch, and design the cross-scale positioning task. Due to the abundance of H&E data, this task can be applied as a normal SSL method to pretrain a model.
In this task, the CL framework is used to bridge the representations of a patch from a local view and a global one, and the pretext token within the region network is supplied with downscaled patches. It is easy to obtain the local understanding of a patch directly from the output features of patch networks. It is yet harder to extract a global understanding of a patch from its surrounding region. Hence, we develop the pretext token into positioning weights over mini-patches from regions, with the help of an explicit positioner as the task raiser, and use the weights to extract the target global features.
@@
Figure 2: Detailed networks of two pretext tasks. (a) Three-branch cross-scale positioning network. (b) Two-branch cross-stain transferring network and the transferer module.
Patch in a local view. The two patch networks constitute a traditional CL framework, i.e., MoCo v3, to provide a basic understanding of patches. The input is a differently augmented view of the patch for each branch. The backbone encoder is denoted as $F(\cdot)$. The outputs of the patch network’s backbone are believed to represent the patch in a local view, as [$\mathbf{y};y^{p}]=F\left([\mathbf{x};\epsilon]\right)$. Then, a typical projector $G(\cdot)$ and predictor $H(\cdot)$ are used to generate the key $\mathbf{k}=G\left(\mathbf{y}\right)$ and query $\mathbf{q}=H\left(\mathbf{k}\right)$ as defined in MoCo v3. The contrastive loss is regarded as the loss in a local view $\mathcal{L}_{\text{local}}$, and is defined as follows:
$\mathcal{L}_{\text{local}}=-\log\frac{\exp{(\mathbf{q}\cdot\mathbf{k}_{+}/\tau)}}{\sum_{i=1}^{B}\exp{(\mathbf{q}\cdot\mathbf{k}_{i}/\tau)}}$ (1)
where $\mathbf{k}_{+}$ is the key of a positive sample (that from a different view of the same image), $B$ the batch size, and $\tau$ the temperature hyper-parameter. The parameters are $\theta$ for the online branch and $\xi$ for the target. $\theta$ is updated with gradient, and a momentum update mechanism (He et al., 2020) is applied $\xi\rightarrow m\xi+(1-m)\theta$, where $m$ is a momentum factor.
@@
$\mathbf{p}=[p_{1},\dots,p_{L}]=\text{Softmax}\left(P\left(\dot{y}^{p}\right)\right)$ (2)
@@
$\mathcal{L}_{\text{global}}=D\left(\mathbf{\hat{q}},\text{stopgrad}\left(\mathbf{k}\right)\right)+D\left(\mathbf{q},\text{stopgrad}\left(\mathbf{\hat{k}}\right)\right)$ (3)
where $D(\mathbf{q},\mathbf{k})=-\frac{\mathbf{q}}{\|\mathbf{q}\|_{2}}\cdot\frac{\mathbf{k}}{\|\mathbf{k}\|_{2}}$ is cosine similarity, and $\text{stopgrad}(\cdot)$ means a gradient stop manipulation (Chen and He, 2021). The local and global losses are then optimized equally.
@@ 3.3 Cross-stain transferring
As mentioned earlier, pathologists cannot always give a detailed diagnosis if only H&E slides are available. They sometimes require more evidence like IHC slides to define the disease. However, limited by the inadequacy of IHC slides, they should detect abnormal regions in H&E slides, find the corresponding regions in IHC slides, and finally derive the expected information. The underlying concept is they obtain structural information from H&E images and further diagnostic information from corresponding IHC regions. Inspired by this, we formulate the relation as paired H&E and IHC images and design the cross-stain transferring task. Considering the scarcity of desired datasets, and a priori structural knowledge in the H&E pretrained model, this task is intended as a quick transferring method from existing H&E foundation models to foundation models of IHC other than a pretraining method.
In this task, the CL framework is used to bridge the representations of a real IHC patch and a pseudo one transferred from the corresponding H&E patch, and the pretext token in the H&E network is furnished with cropped IHC images to provide sufficient stain information. We borrow the concept of adaptive instance normalization (AdaIN) (Huang and Belongie, 2017) that the style of images is rooted in some feature statistics like channel mean and variance, and utilize it to design a transferer as task raiser. The transferer replaces the mean and standard error of H&E features with those of a reference IHC crop, thereby transferring H&E features into IHC styles. The normalized features are viewed as pseudo features of the corresponding IHC patch with injected existing H&E semantics, e.g., separation of structurally normal and abnormal patches. Meanwhile, a basic CL method with negative samples ensures learning the innateness of IHC images.
@@
$=\frac{\mathbf{\dot{y}}-\mu_{\mathbf{\dot{y}}}}{\sigma_{\mathbf{\dot{y}}}}\cdot\sigma_{\dot{y}^{p}}+\mu_{\dot{y}^{p}}$
@@
$\mathcal{L}_{\text{IHC}}=-\log\frac{\exp{(\mathbf{q}\cdot\mathbf{\tilde{k}}_{+}/\tau)}}{\sum_{i=1}^{B}\exp{(\mathbf{q}\cdot\mathbf{\tilde{k}}_{i}/\tau)}}$ (5)
@@ 4 Experiments
These results are based on quantitative experiments with other models. The models first include models using ImageNet (Deng et al., 2009) as baselines, i.e., fully-supervised ViT (Dosovitskiy et al., 2020) denoted as ImageSup and self-supervised ViT via MoCo v3 (Chen et al., 2021) as ImageSSL. Besides, some pathological models are also taken into consideration. Ciga et al. (2022) utilize SimCLR strategy to pretrain a ResNet-18 with large amounts of public pathological data (SimCLR-ciga). Wang et al. provide both a pretrained ResNet-50, RetCCL (Wang et al., 2023a), and a pretrained hybrid architecture of CNN and Swin Transformer, CTransPath (Wang et al., 2022), via customized methods and pathological datasets. These five open-source models are compared throughout the experiments in this section. Pretrained models with ultra-large scale amounts of pathological data, like UNI (Chen et al., 2023) and Virchow (Vorontsov et al., 2023), are postponed to Section 5.3 for a simple discussion since they are not publicly available in this time.
@@ 4.1 Pretraining stage
Datasets. The H&E dataset originates from TCGA  [^1] https://portal.gdc.cancer.gov/, a large-scale public dataset containing genome, epigenome, transcriptome, and image data. In this work, we collect about 30 thousand WSIs from it and select about 11 thousand diagnostic formalin-fixed paraffin-embedded (FFPE) WSIs for training. The patches are cropped under the highest magnification level with a size of $256\times 256$ pixels, and the regions under the second highest level with a size of $1024\times 1024$ pixels. Hence, the physical ratio of size between regions and patches is typically around 8. Finally, 1,623,258 regions and 13,166,437 patches are acquired. The cross-stain dataset originates from HyReCo (Lotz et al., 2022; van der Laak et al., 2021) and BCI dataset (Liu et al., 2022). HyReCo dataset consists of nine datasets of consecutive sections, each containing four slides stained with H&E, CD8, CD45, and Ki67, respectively. Additional PHH3-stained slides are re-stained from the bleached H&E slides. In total, 2,771 pairs of H&E and one stain of IHC are acquired. BCI dataset contains 4,873 pairs of H&E and HER2 images, 3,896 pairs for train, and 977 for test. We only use the 3,896 training pairs. To obtain more training data, we crop these images under another resolution, and finally, 21,126 pairs are used in the second task.
@@
Settings. In the training stage, we use a typical ViT-B/16 backbone and activate automatic mixed precision by PyTorch. For later consistency with the cross-scale positioning task, we use the “avg" mode instead of “token”. Also, considering the gap between ImageNet and our histopathological dataset, we ignore the normalization with ImageNet’s mean and variance. Following MoCo v3, the projector is a three-layer perceptron, and the predictor a two-layer perceptron with a hidden dimension of $4096$. The data augmentation strategies include random cropping and scaling, Gaussian blur, color distortion, random flipping, following BYOL. The batch size is $2048$, $\tau$ is set as $0.2$, and the learning rate is initially $0.00015$ with a scaler of $BatchSize/256$ and updated using a cosine decay schedule with a long warm-up of $40$ epochs. AdamW is used as the optimizer with $0.1$ weight decay.
@@ 4.2 Downstream tasks with H&E images
Cancer tissue identification is one of the most important tasks in computational pathology, especially with H&E images. One formulation of this task is patch-level classification, or supervised image classification, which presents a relatively simple way to test models’ basic understanding of H&E patches. To both study the native ability of models and simulate the real scenarios, a linear probing strategy (Chen et al., 2020) and a full finetuning one are implemented to compare models’ performance.
Another one is WSI-level classification, or weakly-supervised classification, which closely resembles real-world scenarios and is more persuasive in demonstrating models’ global understanding of pathological slides. This task is usually implemented in a multiple instance learning way, and we utilize the attention-based CLAM (Lu et al., 2021) as the framework to perform the MIL process.
@@ 4.2.1 Patch-level tissue subtyping
Datasets. We use the NCT-CRC-HE dataset (Kather et al., 2018) for the patch classification task. It is a collection of histopathology images specifically focused on CRC. It consists of 9 categories of tissue types, with one category representing normal tissues (NORM) and the remaining 8 categories representing colorectal cancer tissues, including adipose (ADI), background (BACK), debris (DEB), lymphocytes (LYM), mucus (MUC), smooth muscle (MUS), cancer-associated stroma (STR) and colorectal adenocarcinoma epithelium (TUM). A total number of 100,000 image patches with a size of $224\times 224$ pixels at $0.5$ microns per pixel (mpp) are used as the training set, while 7,180 image patches are used as the testing set. All training patches are cropped from 86 H&\&&E WSIs and testing patches from 50 H&\&&E WSIs.
@@
Methods CAMELYON16 TCGA-NSCLC TCGA-RCC
@@
Results. In Table 1, we evaluate our H&E model using the linear probing method under different amounts of data. From the result, we can see that our model has shown its superiority across various amounts of training data over other pretrained models. Meanwhile, it can be observed that a generally consistent increasing trend exists with the growth of amounts of training data, but the difference is relatively small for most models. A further study is conducted in Section 5.2 on the training data requirements of foundation models. In Table 2, we present the evaluation of models’ performance under different training strategies using the whole NCT-CRC-HE dataset. The results demonstrate that the proposed model is a good interpreter of H&E images under both a quick linear transferring manner and a thorough full fine-tuning protocol. The performance gain can be owed to the cross-scale positioning task, which enhances the model’s understanding under a broader view. To verify the assumption, an ablating study is discussed in Section 5.1.
@@ 4.2.2 WSI-level classification
Recent works (Campanella et al., 2019; Shao et al., 2021; Lu et al., 2021) have demonstrated the effectiveness of multiple instance learning (MIL) on weakly-supervised classification of WSIs. These MIL-based methods typically follow a two-step approach. First, WSIs are divided into smaller patches to generate patch-level features by utilizing a pretrained model. Second, to generate a slide-level prediction, patch-level features are aggregated using various feature fusion techniques, including recurrent neural network (RNN) or Transformer-based aggregators (Campanella et al., 2019; Shao et al., 2021) and attention-based pooling (Lu et al., 2021).
@@
For pretrained models, we freeze the parameters and utilize the attention-based CLAM as the framework to perform the MIL process. We use Acc and area under the receiver operating characteristic curve (AUC) score to evaluate the WSI classification task. The AUC is calculated using the macro-averaged one when the number of classes is larger than 2. Besides the aforementioned methods, we further include the original CLAM (Lu et al., 2021) and TransMIL (Shao et al., 2021) as baselines. It is worth noting that except for TransMIL, other results are reproductive ones considering the update of datasets.
@@
Datasets. We evaluate this task on three WSI-level datasets: CAMELYON16, TCGA non-small cell lung cancer (NSCLC), and TCGA renal cell carcinoma (RCC). CAMELYON16 dataset was released as part of the Camelyon16 challenge (Bejnordi et al., 2017), which focused on two types of breast cancer classification: benign tissue and metastatic breast cancer. The dataset consists of a total of 399 Whole Slide Images, with 270 WSIs for training and 129 for testing. Although the dataset does provide detailed pixel-level annotations, in the context of weakly-supervised classification, we only utilize the global slide-level annotations, i.e., whether a WSI contains tumor areas or not. TCGA-NSCLC dataset is derived from the TCGA database for two class subtyping: lung squamous cell carcinoma (TCGA-LUSC) and lung adenocarcinoma (TCGA-LUAD). It contains a total of 1053 diagnostic WSIs, including 512 LUSC slides from 478 cases and 541 LUAD slides from 478 cases. TCGA-RCC dataset is another subset of TCGA which includes three subtypes of kidney tumor: kidney chromophobe renal cell carcinoma (TCGA-KICH), kidney renal clear cell carcinoma (TCGA-KIRC), and kidney renal papillary cell carcinoma (TCGA-KIRP). It contains a total of 940 diagnostic WSIs, including 121 KICH slides from 109 cases, 519 KIRC slides from 513 cases, and 300 KIRP slides from 276 cases.
@@
Results. In Table 3, various methods are compared using three different public datasets. Our model shows impressive performance over these three datasets, e.g., +3.1% accuracy gain in CAMELYON16, +0.3% in TCGA-RSCLC, and +1.1% in TCGA-RCC compared with the second best performance. As for other models, pathological models surpass original CLAM and ImageNet models in most cases, but the superiority is not consistent. This might be credited to the strong generalizing ability and global understanding gained from ImageNet, but the results of our models demonstrate that it is effective to use cross-scale positioning tasks to enhance the global understanding of pathological images.

### v2

@@ 3.2 Cross-scale positioning
If one observes how pathologists view slides, the most frequent action of them is likely to be zooming in and out. This phenomenon originates from the diagnostic process wherein pathologists first detect suspicious regions from a global view, then zoom in and define the region with its local surroundings, and subsequently zoom out to inspect the next region of interest. The underlying concept is that they obtain a global and coarse understanding of a WSI under low magnification, while a local but fine understanding of critical regions under high magnification. Inspired by this, we abstract the concept as a larger region containing a smaller patch, and design the cross-scale positioning task. Due to the abundance of H&E data, this task can be applied as a normal SSL method to pretrain a model, resulting in a three-branch architecture shown in Fig. 3 (a). Two of the branches exactly form the original CL framework to provide basic understanding, while the rest one performs the cross-scale positioning task to enhance understanding of different scales.
In this task, the goal is to bridge the representations of a patch from a local view and a global one. The local understanding of a patch is directly obtained from the output features of patch networks. To extract a global understanding, we jointly input patches (as pretext token) and their surrounding regions (as main input) into the region network. We instantiate the task raiser as an explicit positioner module to get positioning weights over mini-patches of regions, and use the weights to extract the desired global features.
@@
Fig. 3: Detailed networks of two pretext tasks. The flow of pretext token is represented by the black arrows, $\epsilon$ is the placeholder, and the task raisers (positioner and transferer) are presented in the white blocks. (a) Three-branch cross-scale positioning network. (b) Two-branch cross-stain transferring network and the transferer module.
Patch in a local view. The two patch networks constitute a traditional CL framework, i.e., MoCo v3, to provide a basic understanding of patches. The input is a differently augmented view of the patch for each branch. The backbone encoder is denoted as $F(\cdot)$. The outputs of the patch network’s backbone are believed to represent the patch in a local view, as [$\mathbf{y};y^{p}]=F\left([\mathbf{x};\epsilon]\right)$. Then, a typical projector $G(\cdot)$ and predictor $H(\cdot)$ are used to generate the key $\mathbf{k}=G\left(\mathbf{y}\right)$ and query $\mathbf{q}=H\left(\mathbf{k}\right)$ as defined in MoCo v3. The contrastive loss is regarded as the loss in a local view $\mathcal{L}_{\text{local}}$, and is defined as,
$\mathcal{L}_{\text{local}}=-\log\frac{\exp{(\mathbf{q}\cdot\mathbf{k}_{+}/\tau)}}{\sum_{i=1}^{B}\exp{(\mathbf{q}\cdot\mathbf{k}_{i}/\tau)}},$ (1)
where $\mathbf{k}_{+}$ is the key of a positive sample (that from a different view of the same image), $B$ the batch size, and $\tau$ the temperature hyper-parameter. The parameters are $\theta$ for the online branch and $\xi$ for the target. $\theta$ is updated with gradient, and a momentum update mechanism [18] is applied as $\xi\rightarrow m\xi+(1-m)\theta$, where $m$ is a momentum factor.
@@
$\mathbf{p}=[p_{1},\dots,p_{L}]=\text{Softmax}\left(P\left(\dot{y}^{p}\right)\right),$ (2)
@@
$\mathcal{L}_{\text{global}}=D\left(\mathbf{\hat{q}},\text{stopgrad}\left(\mathbf{k}\right)\right)+D\left(\mathbf{q},\text{stopgrad}\left(\mathbf{\hat{k}}\right)\right),$ (3)
where $D(\mathbf{q},\mathbf{k})=-\frac{\mathbf{q}}{\|\mathbf{q}\|_{2}}\cdot\frac{\mathbf{k}}{\|\mathbf{k}\|_{2}}$ is cosine similarity, and $\text{stopgrad}(\cdot)$ means a gradient stop manipulation [8]. The local and global losses are then optimized equally.
@@ 3.3 Cross-stain transferring
As mentioned earlier, pathologists cannot always give a detailed diagnosis if only H&E slides are available. They sometimes require more evidence like IHC slides to define the disease. However, limited by the inadequacy of IHC slides, they should detect abnormal regions in H&E slides, find the corresponding regions in IHC slides, and finally derive the expected information. The underlying concept is they obtain structural information from H&E images and further diagnostic information from corresponding IHC regions. Inspired by this, we formulate the relation as paired H&E and IHC images and design the cross-stain transferring task. Considering the scarcity of desired datasets, and a priori structural knowledge in the H&E pretrained model, this task is implemented in a two-branch way without a traditional CL branch and initiated from existing H&E models. The whole process can be seen as we first acquire basic knowledge about pathological images using traditional CL method, then we enhance the global understanding using cross-scale positioning task, and now we use another cross-stain transferring task to explicitly translate H&E understanding into IHC.
In this task, the goal is to bridge the representations of a real IHC patch and a pseudo one transferred from the corresponding H&E patch. The real representations are drawn directly from the IHC network. As for the pseudo ones, we borrow the concept of adaptive instance normalization (AdaIN) [19] that the style of images is rooted in some feature statistics like channel mean and variance, and utilize it to design a transferer as task raiser in the H&E network. After we jointly input IHC references (as pretext token) and their paired H&E patches (as main input) into the H&E network, the transferer replaces the mean and standard error of H&E features with those of IHC reference, thereby transferring H&E features into IHC styles. The normalized features are viewed as pseudo features of the corresponding IHC patch with injected existing H&E semantics, e.g., separation of structurally normal and abnormal patches. Meanwhile, a basic CL method with negative samples ensures learning the innateness of IHC images.
@@
$=\frac{\mathbf{\dot{y}}-\mu_{\mathbf{\dot{y}}}}{\sigma_{\mathbf{\dot{y}}}}\cdot\sigma_{\dot{y}^{p}}+\mu_{\dot{y}^{p}},$
@@
$\mathcal{L}_{\text{IHC}}=-\log\frac{\exp{(\mathbf{q}\cdot\mathbf{\tilde{k}}_{+}/\tau)}}{\sum_{i=1}^{B}\exp{(\mathbf{q}\cdot\mathbf{\tilde{k}}_{i}/\tau)}},$ (5)
@@ 4 Experiments
These results are based on quantitative experiments with other models. The models first include models using ImageNet [11] as baselines, i.e., fully-supervised ViT [14] denoted as ImageSup and self-supervised ViT via MoCo v3 [9] as ImageSSL. Besides, some pathological models are also taken into consideration. Ciga et al. [10] utilize SimCLR strategy to pretrain a ResNet-18 with large amounts of public pathological data (SimCLR-ciga). Wang et al. provide both a pretrained ResNet-50, RetCCL [52], and a pretrained hybrid architecture of CNN and Swin Transformer, CTransPath [53], via customized methods and pathological datasets. The last is UNI [6], a ViT-Large pretrained with over 100 thousand slides using DINO v2 framework. These open-source models are compared throughout the experiments in this section. Pretrained model with ultra-large scale amounts of pathological data, like Virchow [49], is postponed to Section 5.3 for a simple discussion since it is not publicly available at this time.
@@ 4.1 Pretraining stage
Datasets. The H&E dataset originates from TCGA  [^2] https://portal.gdc.cancer.gov/, a large-scale public dataset containing genome, epigenome, transcriptome, and image data. In this work, we collect about 30 thousand WSIs from it and select about 11 thousand diagnostic formalin-fixed paraffin-embedded (FFPE) WSIs for training. The patches are cropped under the highest magnification level with a size of $256\times 256$ pixels, and the regions under the second highest level with a size of $1024\times 1024$ pixels. Hence, the physical ratio of size between regions and patches is typically around 8. Finally, 1,623,258 regions and 13,166,437 patches are acquired. The cross-stain dataset originates from HyReCo [35, 27] and BCI dataset [34]. HyReCo dataset consists of nine datasets of consecutive sections, each containing four slides stained with H&E, CD8, CD45, and Ki67, respectively. Additional PHH3-stained slides are re-stained from the bleached H&E slides. In total, 2,771 pairs of H&E and one stain of IHC are acquired. BCI dataset contains 4,873 pairs of H&E and HER2 images, 3,896 pairs for train, and 977 for test. We only use the 3,896 training pairs. To obtain more training data, we crop these images under another resolution, and finally, 21,126 pairs are used in the second task.
@@
Settings. In the training stage, we use a typical ViT-B/16 backbone and activate automatic mixed precision by PyTorch. For later consistency with the cross-scale positioning task, we use the “avg” mode instead of “token”. Also, considering the gap between ImageNet and our histopathological dataset, we ignore the normalization with ImageNet’s mean and variance. Following MoCo v3, the projector is a three-layer perceptron, and the predictor a two-layer perceptron with a hidden dimension of $4096$. The data augmentation strategies include random cropping and scaling, Gaussian blur, color distortion, random flipping, following BYOL. The batch size is $2048$, $\tau$ is set as $0.2$, and the learning rate is initially $0.00015$ with a scaler of $\text{BatchSize}/256$ and updated using a cosine decay schedule with a long warm-up of $40$ epochs. AdamW is used as the optimizer with $0.1$ weight decay.
@@ 4.2 Downstream tasks with H&E images
Cancer tissue identification is one of the most important tasks in computational pathology, especially with H&E images. One formulation of this task is patch-level classification, or supervised image classification, which presents a relatively simple way to test models’ basic understanding of H&E patches. To both study the native ability of models and simulate the real scenarios, a linear probing strategy [7] and a full finetuning one are implemented to compare models’ performance.
Another one is WSI-level classification, or weakly-supervised classification, which closely resembles real-world scenarios and is more persuasive in demonstrating models’ global understanding of pathological slides. This task is usually implemented in a multiple instance learning way, and we utilize the attention-based CLAM [37] as the framework to perform the MIL process.
@@ 4.2.1 Patch-level tissue subtyping
Datasets. We use the NCT-CRC-HE dataset [24] for the patch classification task. It is a collection of histopathology images specifically focused on CRC. It consists of 9 categories of tissue types, with one category representing normal tissues (NORM) and the remaining 8 categories representing colorectal cancer tissues, including adipose (ADI), background (BACK), debris (DEB), lymphocytes (LYM), mucus (MUC), smooth muscle (MUS), cancer-associated stroma (STR) and colorectal adenocarcinoma epithelium (TUM). A total number of 100,000 image patches with a size of $224\times 224$ pixels at $0.5$ microns per pixel (mpp) are used as the training set, while 7,180 image patches are used as the testing set. All training patches are cropped from 86 H&\&&E WSIs and testing patches from 50 H&\&&E WSIs.
@@
Methods CAM16 NSCLC RCC
@@
Settings. For linear evaluation on the NCT-CRC-HE dataset, following CTransPath, we use Adam as the optimizer with a batch size of $96$. The initial learning rate is set to $0.0003$. Data augmentations of random horizontal, vertical, and 90-degree flips and random cropping are used. For the full fine-tuning, the learning rate for the linear classifier is set to $0.0003$, while the learning rate for the rest part of the pretrained model is set to $0.00003$. The maximum number of epochs is set to $50$ for all models to converge.
Results. In Table 1, we evaluate our H&E model using the linear probing method under different amounts of data. From the result, we can see that our model performs well across various amounts of training data over other pretrained models. Meanwhile, it can be observed that a generally consistent increasing trend exists with the growth of amounts of training data, but the difference is relatively small for most models. A further study is conducted in Section 5.2 on the training data requirements of foundation models. Notably, the giant UNI shows a dominant performance when the training data is extremely limited, which demonstrates its general interpretability of pathological images. In Table 2, we present the evaluation of models’ performance under different training strategies using the whole NCT-CRC-HE dataset. The results demonstrate that the proposed model is a good interpreter of H&E images under both a quick linear transferring manner and a thorough full fine-tuning protocol. The performance gain can be owed to the cross-scale positioning task, which enhances the model’s understanding under a broader view. To verify the assumption, an ablating study is discussed in Section 5.1. UNI also provides decent performance, which shows its great understanding in pathology and powerful ViT-Large architecture.
@@ 4.2.2 WSI-level classification
Recent works [4, 45, 37] have demonstrated the effectiveness of multiple instance learning (MIL) on weakly-supervised classification of WSIs. These MIL-based methods typically follow a two-step approach. First, WSIs are divided into smaller patches to generate patch-level features by utilizing a pretrained model. Second, to generate a slide-level prediction, patch-level features are aggregated using various feature fusion techniques, including recurrent neural network (RNN) or Transformer-based aggregators [4, 45] and attention-based pooling [37].
@@
For pretrained models, we freeze the parameters and utilize the attention-based CLAM as the framework to perform the MIL process. We use Acc and area under the receiver operating characteristic curve (AUC) score to evaluate the WSI classification task. The AUC is calculated using the macro-averaged one when the number of classes is larger than 2. Besides the aforementioned methods, we further include the original CLAM [37] and TransMIL [45] as baselines. It is worth noting that except for TransMIL, other results are reproductive ones considering the update of datasets.
@@
Datasets. We evaluate this task on three WSI-level datasets: CAMELYON16 (CAM16), TCGA non-small cell lung cancer (NSCLC), and TCGA renal cell carcinoma (RCC). CAMELYON16 dataset was released as part of the Camelyon16 challenge [3], which focused on two types of breast cancer classification: benign tissue and metastatic breast cancer. The dataset consists of a total of 399 Whole Slide Images, with 270 WSIs for training and 129 for testing. Although the dataset does provide detailed pixel-level annotations, in the context of weakly-supervised classification, we only utilize the global slide-level annotations, i.e., whether a WSI contains tumor areas or not. TCGA-NSCLC dataset is derived from the TCGA database for two class subtyping: lung squamous cell carcinoma (TCGA-LUSC) and lung adenocarcinoma (TCGA-LUAD). It contains a total of 1053 diagnostic WSIs, including 512 LUSC slides from 478 cases and 541 LUAD slides from 478 cases. TCGA-RCC dataset is another subset of TCGA which includes three subtypes of kidney tumor: kidney chromophobe renal cell carcinoma (TCGA-KICH), kidney renal clear cell carcinoma (TCGA-KIRC), and kidney renal papillary cell carcinoma (TCGA-KIRP). It contains a total of 940 diagnostic WSIs, including 121 KICH slides from 109 cases, 519 KIRC slides from 513 cases, and 300 KIRP slides from 276 cases.
Settings. For the weakly-supervised WSI classification task, following CLAM, we freeze our pretrained model and use the Adam as the optimizer with a batch size of 1 (WSI/bag) and a weight decay of $0.00001$. The learning rate is set to $0.0002$, and the epochs to $50$. For the CAMELYON16 dataset, we adopt the official data split in the CAMELYON16 challenge. For TCGA-NSCLC and TCGA-RCC datasets, we use 5-fold Monte Carlo cross-validation to obtain a more stable result. Each WSI is cropped into non-overlapping small patches after removing the background area with a filter of saturation less than 15 in the CAMELYON16 dataset, and 8 in the TCGA dataset.
@@
Results. In Table 3, various methods are compared using three different public datasets. UNI achieves all best across these three datasets, which demonstrate its power in understanding pathological slides. Excluding UNI, our model shows great performance over these three datasets, e.g., +3.1% accuracy gain in CAMELYON16, +0.3% in TCGA-NSCLC, and +1.1% in TCGA-RCC. This demonstrate that it is effective to use cross-scale positioning tasks to enhance the global understanding of pathological images. As for other models, pathological models surpass original CLAM and ImageNet models in most cases, but the superiority is not consistent. This might be credited to the strong generalizing ability and global understanding gained from ImageNet, especially when the visual encoder is frozen.